OpenStudy (anonymous):

Find all t in the interval [0, 2π] satisfying (cos t)2 − 5 cos t + 4 = 0.

OpenStudy (anonymous):

\[\text{Let } x = \cos t\]\[\cos^2(t) - 5\cos t + 4 = 0\]\[\implies x^2 - 5x + 4 = 0\] use the quadratic formula..

OpenStudy (anonymous):

Then sub your results back to find the value of t.

OpenStudy (anonymous):

Thank youuu!!!

OpenStudy (anonymous):

i got the question wrong, do you have an answer by chance?